OpenStudy (anonymous):

Period 1 has 15 students in it and a test average of 86%. Period 2 has 21 students in it and an average of 88%. Period 3 has 12 students in it and an average of 95%. Use weighted means to find the overall average of the classes. 87.3% 88.7% 89.1% 91%

OpenStudy (kirbykirby):

\(15+21+12=48\) So:\[ \frac{15}{48}(0.86)+\frac{21}{48}(0.88)+\frac{12}{48}(0.95)=?\] and convert the decimal into percent again